OpenStudy (anonymous):

the altitude of an object in feet is given by h=16t^2+8t+48. What is the maximum altitude ?After how many seconds "t" will it strike the ground ?

OpenStudy (amistre64):

according to that equation; itll never hit the ground

OpenStudy (anonymous):

then what is its maximum altitude ?

OpenStudy (amistre64):

according to your equation; gravity will continue to push it out into infinity

OpenStudy (anonymous):

hmmmm

OpenStudy (amistre64):

you sure it aint -16t^2?

OpenStudy (anonymous):

ohh lol IT IS . i read the question wrong sorry

OpenStudy (amistre64):

:) -16t^2 is the result of gravity pulling the object back to earth ...

OpenStudy (amistre64):

well, there are a few ways we can go about this; which ones do you know of?

OpenStudy (anonymous):

none

OpenStudy (anonymous):

can u please explain

OpenStudy (amistre64):

when we plot this on a graph; we get a shape that looks like a hill, its called a parabola

OpenStudy (amistre64):

the top of that hill is the highest point it reaches and it reaches that point in about half the time

OpenStudy (amistre64):

since this started at a height of 48; that measurement will be off but we can account for it

OpenStudy (amistre64):

the most basic way i know of to find the high point, also known as a vertex, is to determine what make the equation equal 0, so we factor it

OpenStudy (amistre64):

-16t^2+8t+48 = 0 ; we can factor out a common -8 first of all -8(2t^2 -t -6) = 0; that -8 is useless to us since -8(0) = 0. so we can ignore it 2t^2 -t -6 = 0 ; my method employs multiplying the 1st and 3rd #s 2(-6) = -12 ; now find factors of -12 that add up to -1 (our 2nd#) -4*3 = -12 ; and -4+3 = -1 this gives us the factors of our equation to be: (t-4/2) (t+3/2) = 0 ; simplify and if its a fraction left over stick the bottom in front (t-2) (2t+3) = 0 when t=2 or t=-3/2 lets plot these on a number line and find the middle: t <...........-3/2..........(C)............2..........> 1/4 when t = 1/4 we are at our highest point; and when t=2 we are back on the ground

OpenStudy (amistre64):

another way to find the middle is to add them together and divide it in half: 2-1.5 = .5 ; .5/2 = .25 to find the height, just use .25 in the equation in the place of t and calculate